### v1

# Verbalized Algorithms

Abstract
Instead of querying LLMs in a one-shot manner and hoping to get the right answer for a reasoning task, we propose a paradigm we call verbalized algorithms (VAs), which leverage classical algorithms with established theoretical understanding. VAs decompose a task into simple elementary operations on natural language strings that they should be able to answer reliably, and limit the scope of LLMs to only those simple tasks. For example, for sorting a series of natural language strings, verbalized sorting uses an LLM as a binary comparison oracle in a known and well-analyzed sorting algorithm (e.g., bitonic sorting network). We demonstrate the effectiveness of this approach on sorting and clustering tasks.
1 Introduction
Imagine receiving thousands of customer complaints and having to quickly sort them by how urgently they need support. Feeding such a large set of input strings blindly into an LLM is ineffective and impractical, both due to its limited context length and its limited ability to perform sorting tasks. LLMs provide no guarantees on the correctness of their outputs for computational reasoning tasks, such as sorting, beyond anecdotal, empirical observations that they sometimes do.
In contrast, traditional disciplines of computer science often approach such a task using task decomposition. For example, a merge sort decomposes the entire sorting task into several sorting tasks over individual sublists and combines the results, where, at the limit, there is an elementary binary comparison operator that compares individual pairs of objects in the list. These traditional algorithms, given an accurate oracle, guarantee the correctness of their outputs and the average/worst-case runtime.
In this work, we propose an abstract framework called “verbalized algorithms”. Given an informal representation of a formal task, a verbalized algorithm operates by replacing some basic atomic operations in a classical algorithm with LLM-driven text-based ones while retaining the high-level control flow of the initial algorithm, thereby largely retaining the theoretical guarantees of the original algorithm.
This has advantages over another common approach to similar tasks that we call the “formalization-based approach”. Leveraging the machine translation ability of LLMs, it converts an informal task representation into a formal, symbolic representation, solves it with a specialized combinatorial solver, and then converts the solution back to the informal, natural language representation for the user. Existing work based on this approach includes LLM-based generation of PDDL (McDermott, 2000) for classical planning (Xie, 2020; Guan et al., 2023; Li et al., 2024; Fine-Morris et al., 2024; Oswald et al., 2024), Prover9 input language for theorem proving (Olausson et al., 2023), or SMT solver input for optimization (Hao et al., 2024). The issue with this approach is twofold: First, it requires grammar-following capability for a complex target formal language. Second, while the downstream solver requires a semantically and logically accurate input, the vagueness and informality in the natural language input often cause LLMs to generate inaccurate and inconsistent translations.
2 Verbalized Algorithms
While whether LLMs can reason or not is up for both philosophical and scientific debates, it is more widely accepted that they can solve some much simpler tasks with high fidelity, such as text-based classification and question-answering. Assuming those capabilities, we propose verbalized algorithms (VAs), an LLM-based reasoning strategy that limits the scope of the LLM’s usage to simple tasks.
Definition 1 (Verbalized Algorithms).
Let $T[\tau]$ be a computational task (e.g., sorting) over objects of type $\tau$ and let $A[\tau]$ be an algorithm that solves $T[\tau]$ with an $n$-arg boolean oracle $f[\tau]:\tau,\ldots,\tau\rightarrow{\left\{\top,\bot\right\}}$. A Verbalized Algorithm instantiates $A[\tau]$ with $\tau=\texttt{str}$ by implementing $f[\texttt{str}]$ as a query to an LLM whose output is constrained to $\left\{\text{``yes''},\text{``no''}\right\}$} which maps to $\left\{\top,\bot\right\}$}.
VAs have three categories: Naive VAs, which treat the LLM’s answer as ground truth, Robust VAs, which treat it as a noisy oracle with a certain error rate and quantify the rate of error in the final output, and Probabilistic VAs, which use the probability/logits of LLM outputs to explicitly quantify the uncertainty of the result and return a distribution of outputs. In this workshop paper, we focus on the naive and the robust variants due to space, leaving the probabilistic variants for future work.
2.1 Verbalized Sorting
Verbalized Sorting sorts a list of natural language strings $X$ using LLM’s question answering capability as an atomic comparison oracle. Each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query such as “Is X better than Y? X: $x_{1}$ Y: $x_{2}$”, where the next output token is restricted to “yes” and “no”. The model output is mapped to $\top/\bot$ if the sampled output is “yes” / “no”, respectively. We evaluated the accuracy of several non-VA and VA approaches. Method, dataset, and prompt details are available in the appendix Sec. A. All results are averages of 5 random seeds.
Constraint Decoding Baseline encodes the list $X$ in a markdown format and asks the model to return a sorted list using structured-decoding (Willard and Louf, 2023) to constrain the output to be a list whose elements are guaranteed to be one of the input strings. Due to the lack of a computationally efficient way to constrain the list to (1) be free of duplicates and (2) represent the same set of strings as the input, we (1) remove the duplicates and (2) append missing elements to the end of the list.
Scoring-based approaches ask the model to assign a score between 1.0 and 5.0 (down to one decimal) to each sentence using regex-constrained decoding, then sort the sentences using the scores. It is a form of formalization-based approach (the list of scores represents the formal problem). We tested two variants: I.i.d Scoring scores $x_{i}\in X$ independently, and Autoregressive Scoring scores all strings in a single query, allowing the model to see the scores of other elements.
VA Powersort overrides Python’s __gt__ operator with an LLM call and applies the built-in sorted function, which uses Powersort (Munro and Wild, 2018) that runs $O(n\log n)$ comparisons worst-case, the best runtime complexity achievable by comparison sort. VA Bitonic sorting network uses the Bitonic sorting network (Batcher, 1968) from parallel sorting literature to improve the throughput. It performs $O(n(\log n)^{2})$ comparisons, but runs in $O((\log n)^{2})$ time due to parallelism. VA Robust Powersort/Bitonic Sorting Network makes individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting. For an odd $K$, if the oracle is correct with probability $p$, the voting result is correct with at least probability $1-e^{-K(2p-1)^{2}/2}$ (Emamjomeh-Zadeh and Kempe, 2018) due to Hoeffding bound (Hoeffding, 1963).
We evaluated the approaches above on Amazon reviews benchmark (Hou et al., 2024). We selected 30 products from the “Video Games” category, which, for each integer rating between 1 and 5, have at least 5 non-empty reviews that are voted as “helpful” by other customers. The task is to sort a shuffled list of 25 selected reviews (5 for each rating) in increasing order of positive sentiment. To measure the accuracy, we compute the Kendall-Tau score (Kendall, 1938) between the output list $Y$ and the ground-truth output list $Y^{*}$, which is sorted by the true rating, averaged over the dataset. We used Qwen3-1.7B and Qwen3-32B-AWQ models (Team, 2025) with thinking enabled.
First, Table 1 (1) shows how often the constraint-decoding baseline violates the sorting invariance, i.e., how many elements are duplicated or missing in the output on average. While these violations become rarer as the model gets larger, they do not disappear. This highlights the lack of theoretical guarantees in LLM-based reasoning, which does not occur in VAs and formalization-based approaches.
Next, Table 1 (2) shows the average Kendall-Tau scores. Constraint-decoding baseline performed the worst overall, and while the scoring approaches can work with a large 32B model, their performance dwindles when the model is smaller. Robust VA with 32B achieved the best results. The strength of VA is most pronounced when using the small 1.7B model, where other approaches failed. Notably, VAs with a 1.7B model outperformed the baseline with a 32B model, clearly showing the advantage of using the sorting algorithms developed in the literature.
Finally, Table 2 compares the runtime of each approach. VA Bitonic Sorting Network improves over the VA Powersort, and both VAs can be faster than the constrained-decoding baseline, even with robustification (x3 parallel queries). This shows that VAs can directly benefit from the theoretical results in the parallel sorting literature, in contrast to the recent LLM reasoning literature that tends to detach from existing discrete algorithm studies. Note that the parallel speedup in bitonic sort does not reach the theoretical limit (x11.4; 171 comparators and 15 steps for 25 inputs) because the maximum batch size is limited by the GPU; multi-GPU experiments with data-parallelism would improve this.
Table 1: (1) Average number of sorting invariance violations by the constraint decoding baseline. “Duplicates” indicates the number of duplicates found in the output, and “missing” indicates the number of input elements that are missing in the output, for 25 input strings. (2) Average Kendall-Tau scores. Top-3 are highlighted in bold. The score of 1, 0, and -1 indicates perfect rank correlations, no correlations, and perfect anti-correlations, respectively.
(1) Invariance (2) Accuracy (Kendall-Tau scores) Models Violations Constraint Scoring Naive VA Robust VA ($K=3$) (Qwen3) duplicates missing Decoding i.i.d. AR Bitonic Powersort Bitonic Powersort 1.7B 1.39 9.63 0.00 0.18 0.05 0.30 0.33 0.33 0.44 32B-AWQ 0.49 2.75 0.30 0.57 0.39 0.39 0.56 0.39 0.58
Table 2: Average total elapsed time in seconds for processing all sorting queries. Bitonic sorting network also shows the relative speedup over the sequential sort (Powersort).
Models Scoring Naive VA Robust VA ($K=3$) (Qwen3) Baseline i.i.d. AR Bitonic Powersort Bitonic Powersort 1.7B 7468 630 205 2338 (x2.12) 4975 5060 (x1.45) 7362 32B-AWQ 7280 1430 626 8556 (x2.18) 18679 20350 (x1.30) 26471
2.2 Verbalized Clustering
Verbalized Clustering attempts to cluster a set of strings $S$ into a family of subsets $\mathcal{S}$ based on their similarity. The problem with text-based clustering is that the distance function for natural language strings is difficult to define. All approaches to this task are embedding methods. Method, dataset, and prompt details are available in the appendix Sec. B. All results are averages of 5 random seeds.
Jaccard Baseline creates a distance matrix using Jaccard distance of token sets of strings, then applies sklearn’s MDS (Kruskal, 1964) to create an embedding. Sentence Transformer (ST) Baseline uses Qwen3-8B-Embedding directly, which was selected as the top performer in the MTEB leaderboard (Muennighoff et al., 2022). VA t-STE: This is a VA based on Student-t Stochastic Triplet Embedding (Van Der Maaten and Weinberger, 2012, t-STE), which makes triplet comparison queries to LLM. Triplet comparison takes three elements $x$, $y$ and $z$ in a metric space $\|\cdot\|$ and query if $\|x-y\|<\|x-z\|$. We ask the LLM a question such as “Is X closer to Y than is to Z? X: $x$, Y: $y$, Z: zz”. For each $x\in S$, we randomly sample $k$ queries from $S^{3}$ and their results from an LLM, then generate $D$-dim embeddings by minimizing the number of violations to the results. We show the results with hyperparameters $k\in{\left\{100,200\right\}},D=8$.
We evaluated the methods across four datasets. Medarxiv (Muennighoff et al., 2022) contains the titles and the categories of medical papers on MedArxiv. The category forms a cluster. IT Field Defects consists of $78$ technical IT field defect reports submitted to our internal defect tracking system (unpublished), where some reports are manually marked as duplicates of another, forming a cluster. xkcdcolors-h/l are two datasets of 100 colors sampled from (Munroe, 2010). Each entry has a non-standard color name (e.g., baby blue, vomit) and a corresponding RGB value. The difference between h and l variants is that we defined 6 ground truth clusters with sklearn.AgglomerativeClustering using a lexicographic distance that prioritizes hue and lightness in the HSL color model, respectively. Assuming that colors tend to be more strongly associated with hues, xkcdcolors-l represents an out-of-distribution (OOD) task.
We evaluate the embeddings with a score defined in Eq. 1, which is the average rank of the members of the cluster each string belongs to, when sorted by the Euclidean distance. The score is normalized to $[0,1]$, where 1 indicates that cluster members are embedded closer. Given the input $S$ and a string $s\in S$, let $C(s)$ be the ground-truth cluster such that $s\in C(s)$. Let $i=\textsc{rank}(s,s^{\prime})\in{\left\{0,1,\ldots|S|-1\right\}}$ be an integer such that, when sorted by the pairwise Euclidean distance $d(e(s),e(s^{\prime\prime}))$ between $s$ and a string $s^{\prime\prime}$ across $s^{\prime\prime}\in S$ using embedding $e$, then $d(e(s),e(s^{\prime}))$ ranks $i$-th. The score is defined as:
score $=\textstyle\quad\frac{1}{|S|}\sum_{s\in S}\frac{1}{|C(s)|-1}\sum_{s^{\prime}\in C(s)\setminus{\left\{s\right\}}}1-\frac{\textsc{rank}(s,s^{\prime})}{|S|}\in[0,1]$ (1)
Table 3 shows the results. Jaccard fared surprisingly well in IT Defects, where the input often contains long output logs whose tokens carry significant information, but it underperformed in tasks that require semantic comprehension. Sentence Transformer achieves strong performance in the standard clustering tasks such as Medarxiv and IT Defects. However, it fell short in xkcdcolors-h/l that requires information about the distance criteria. The score 0.53 in the xkcdcolors-l is close to 0.5, i.e., its distance ranking is only marginally better than a dice roll. VA t-STE outperformed traditional baselines with $k=200$. We highlight its strong generalization in OOD tasks where other approaches failed, due to its strong algorithmic bias combined with flexible LLM queries.
Table 3: Clustering scores across methods and benchmarks. The top 2 results are highlighted in bold.
method model $D=$ $k=$ Medarxiv IT Defects xkcdcolors-h xkcdcolors-l Jaccard 8 0.556 0.767 0.601 0.472 ST Qwen3-Emb…d-8B 4096 0.827 0.936 0.718 0.531 triplet Qwen3-4B 8 100 0.831 0.831 0.823 0.704 triplet Qwen3-4B 8 200 0.858 0.895 0.823 0.691 triplet Qwen3-32B-AWQ 8 100 0.842 0.908 0.851 0.748 triplet Qwen3-32B-AWQ 8 200 0.858 0.939 0.850 0.742
3 Conclusion, Related Work, Discussion and Future Work
We presented Verbalized Algorithms (VAs), a disciplined paradigm for performing algorithmic reasoning with LLMs. It benefits from a strong prior in the golden classics of discrete algorithms published half a century ago, significantly improving the reasoning accuracy even with a small model. It also leverages their established theoretical results to improve the runtime of such reasoning.
Despite significant recent findings in reasoning models, which combine RL-based self-training with “think tags” (Guo et al., 2025), they rarely explicitly perform systematic backtracking and search. Whether such systems can reason or not has been hotly debated in the context of LLM-based planning (Valmeekam et al., 2023). Previous work on solving planning tasks using LLM (Liu et al., 2023; Hirsch et al., 2024) demonstrates varying degrees of success and limitations (e.g., failure in mystery due to sensitivity to the vocabulary). Thought-of-Search (Katz et al., 2024) simulates a systematic search algorithm, which could be seen as an instance of VAs.
While this paper explored only a limited number of applications due to space, we are already pursuing several exciting verbalizations, including submodular maximization and pareto-front computation for diverse / non-dominated document retrieval, DPLL for combinatorial reasoning, and median/top-$k$ selection, etc. The potential of this paradigm is large, and we invite various ideas for unique verbalizations and new applications. For example, can Floyd’s tortoise-and-hare cycle detection algorithm for linked lists be verbalized to perform a tautology detection? Garbage collection algorithms to detect unimportant sentences in a paper draft?
VAs require multiple LLM calls by nature, which creates a performance bottleneck. An important avenue of future work includes leveraging efficient prefix caching, e.g., reusing the KV-cache of the prefix “Is X brighter than Y? X: $x_{1}$ Y:”, theoretically analyzing the “prefix-cache-friendliness” of various verbalized algorithms, and deriving the error bounds of VAs using past algorithmic literature on noisy oracles.
Appendix
Appendix A Sorting: Experimental Details
A.1 Bitonic Sorting Network
Bitonic sorting networks (Batcher’s Algorithm) (Batcher, 1968) allow us to optimize the sorting of a given list. The comparisons in a bitonic sorting network are not data-dependent, allowing for more feasible parallelization and speed proportional to depth $D$. Given an input list, a bitonic sorting network originally sorts both halves of the list, subsequently conducts an odd-even mergesort on both halves of the list and finally switches adjacent keys.
A sequence $a_{0},\ldots,a_{n-1}$ of $n$ elements is called bitonic when, for some $i$ (called an inflection point), $a_{0}\leq a_{1}...\leq a_{i}\geq a_{i+1}...\geq a_{n-1}$ (i.e., the first part up to $i$ is monotonically increasing and the second part is monotonically decreasing). Assume $n$ is a power of 2 and a sequence $a_{0},\ldots,a_{n-1}$ is bitonic with inflection point $i=\frac{n}{2}$. Then, we construct two sequences $s_{1}$ and $s_{2}$ with a following compare-and-exchange process:
$s_{1}$ $=\min(a_{0},a_{\frac{n}{2}}),\min(a_{1},a_{\frac{n}{2}+1}),...\min(a_{\frac{n}{2}-1},a_{n-1})$
$s_{2}$ $=\max(a_{0},a_{\frac{n}{2}}),\max(a_{1},a_{\frac{n}{2}+1}),...\max(a_{\frac{n}{2}-1},a_{n-1}).$
$s_{1}$ and $s_{2}$ are both bitonic sequences. Moreover, any element in $s_{2}$ are larger than any element in $s_{1}$, i.e., $\forall x\forall y$ $x\in s_{1}$, $x\in s_{2}$, $x<y$. The recursive compare-and-exchange process above is applied on $s_{1}$ and $s_{2}$ for $\log(n)$ times to sort both halves of the list.
For some input sizes, there are optimal sorting networks which require less comparisons or less parallel steps than Bitonic sorter does (Knuth, 1997). Bitonic sorting network can switch to such a network during the recursion, further increasing the performance.
A.2 Kendall-Tau Score
The Kendall Tau correlation coefficient Kendall (1945) is a non-parameteric statistic that measures how much order is preserved between two sets of rankings. Kendall-Tau score has several ways to compute ($\tau_{a},\tau_{b},\tau_{c}$) depending on the existence of ties or the nature of the observations. Here we describe $\tau_{b}$ we used in our paper.
Assume each observation consists of an input variable $X_{i}$ and an output variable $Y_{i}$. Consider two observations $(X_{i},Y_{i})$ and $(X_{j},Y_{j})$. These observations are called concordant if $X_{i}<X_{j}$ and $Y_{i}<Y_{j}$, or $X_{i}>X_{j}$ and $Y_{i}>Y_{j}$. Conversely, they are called discordant if $X_{i}<X_{j}$ and $Y_{i}>Y_{j}$, or $X_{i}>X_{j}$ and $Y_{i}<Y_{j}$. If $X_{i}=X_{j}$ and/or $Y_{i}=Y_{j}$, the observations are called tied.
Given $n$ observations $O$, enumerate all $n\choose 2$} pairs, which we denote as $N$. NN is then comprised of concordant pairs $P$, discordant pairs $Q$, pairs tied on the $X$ variable denoted as $T$, pairs tied on the $Y$ variable denoted as $U$, as well as pairs tied on both variables denoted as $B$. Kendall $\tau_{b}$ is then defined as:
$\tau_{b}=\frac{|P|-|Q|}{\sqrt{(|P|+|Q|+|T|)*(|P|+|Q|+|U|)}}\in[-1,1].$
A $\tau_{b}$ close to 1-1 indicates a negative correlation between two rankings and A $\tau_{b}$ close to $1$ indicates aligned rankings.
A.3 LLM Query Prompts
All methods uses a common shared criteria for comparison. In the Amazon dataset used in the study, the criteria prompt is:
criteria = ‘‘You will be given two product reviews, X and Y, written by various customers about the same product. Is X more favorable to the product than Y?’’
We used the following query prompts:
Constrained-decoding baseline:
f"Sort the following list in the increasing order of quality that is compared in the following comparison criteria: ’{criteria}’. That is, for a pair X and Y, if the answer to the criteria is ’yes’, X must appear after Y in the output list. "
I.i.d Scoring:
[…]
Autoregressive Scoring:
[…]
Verbalized Sorting:
f"{criteria} Answer with a single word, yes or no. X: ’{self.string}’ Y: ’{other.string}’"
Appendix B Clustering: Experimental Details
B.1 Jaccard distance
This implements a baseline clustering method using lexical similarity. Let $T_{i}$ represent the set of tokens from a string $s_{i}\in S$. For any two strings $s_{i}$ and $s_{j}$ with corresponding token sets $T_{i}$ and $T_{j}$, the Jaccard distance is computed as:
$J(s_{i},s_{j})=1-\frac{|T_{i}\cap T_{j}|}{|T_{i}\cup T_{j}|}$
where $|T_{i}\cap T_{j}|$ represents the cardinality of the intersection and $|T_{i}\cup T_{j}|$ represents the cardinality of the union.
B.2 t-STE
Van Der Maaten and Weinberger (2012) proposed a heavy-tailed kernel to measure local similarities between data points. In particular, we opt to use a Student-$t$ kernel with $\alpha$ degrees of freedom by defining:
$p_{ij\ell}=\frac{\left(1+\frac{\lVert x_{i}-x_{j}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}}{\left(1+\frac{\lVert x_{i}-x_{j}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}+\left(1+\frac{\lVert x_{i}-x_{\ell}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}}$
We refer to the resulting technique as t-Distributed STE (t-STE). Our implementation sets $\alpha=D-1$, where $D$ is the number of dimensions in the resulting embedding.
The formulation of the triplet probabilities is motivated by the success of unsupervised dimensionality reduction technique that also employ heavy-tailed similarity kernels (Maaten and Hinton, 2008).
B.3 LLM Query Prompts
Triplet comparison query uses the following shared prompt, where criteria is a dataset-specific prompt.
[…]
The dataset-specific prompt is as follows:
Medarxiv dataset:
criteria = f"You will be given three paper titles X, Y, and Z. Does the paper Z discuss a topic that is closer to the topic in X than to the topic in Y?"
IT Defect dataset:
criteria = f"You will be given three defect reports X, Y, and Z, sent by the customers of our GPFS filesystem storage system. They may contain stack traces of various programs (python, kernel drivers, ...). Similar repetitive lines are omitted as [...]. Considering the cause of the defects (i.e. ignoring timestamp/IP/MAC addresses), are X and Z more similar than Y and Z are? "
xkcdcolors-h dataset:
criteria = f"Which of the following two colors, X and Y, is more similar to the color Z? Focus on the hue in the HSL model first, then other aspects such as lightness and saturation. For example, Red and Pink are more similar than Red and Blue are. If you are not familiar with the given color name, e.g., a made-up color name like ’puke brown’, try to imagine what the color would look like, e.g., puke would generally have a lighter shade. "
xkcdcolors-l dataset:
criteria = f"Which of the following two colors, X and Y, is more similar to the color Z? Focus on the lightness in the HSL model first, then other aspects such as hue and saturation. For example, Red and Blue (both lightness = 50 If you are not familiar with the given color name, e.g., a made-up color name like ’puke brown’, try to imagine what the color would look like, e.g., puke would generally have a lighter shade. "
B.4 IT Defect Reports Dataset
Let $D$ denote the set of all defects. A subset $D^{\prime}\subseteq D$ was manually labeled as duplicates of several other defects in $D$. We selected 78 data points so that the final dataset contains 25 such clusters, each of size at least 2.
Since these labels are directional and sparse, we define the cluster by making the link undirected and computing the connected components, then treating a set of elements in each connected component as a cluster.
The following is a redacted/simplified defect records that appear in the dataset.
Defect 1
  "id": 347852,
  "description":
  "System login:
    EMS - X.XX.XXX.XXX root/cluster
    S6K w/ Falcon attached - lothal-qa6-[1 && 2] Code levels:
    S6k - 6230_B19
    FCM4 - 4_3_3
    Falcon - 5342 --> 5343
   Snap location:
    glogin:/u/DUMPS/R.347852/
   What I was doing:
    I was upgrading Falcon enclosure from 5342 to 5343.
   What I expected:
    ’XXXXfirmware’ will start and the Falcon will go through the upgrade process.
   What happened:
    ’mmchrimware’ completed but there are errors printed.
    The Falcon does appear to have the updated firmware.
    However, there are issues when trying to get a print
    from the Left module when running ’essenclosures’.
   Analysis:
    sg91:13:/dev/sg91::naa.500093d006585000/10910::[1.0.91.0]:/dev/sg91:XXX:5149-091:5342
    [10:48:31.258608008] XXXXXXFirmwareVer2 checking if enclosure needs to be updated
    [10:48:31.260567945] XXXXXXFirmwareVer2 firmware file found
    [10:48:31.262950549] XXXXXXFirmwareVer2 -> ...
    [10:48:31.350737778] XXXXXXFirmwareVer2 ...
    [10:48:31.362309567] XXXXXXFirmwareVer2 lothal-qa6-1-hs: [E] Error 1 updating firmware ...
    sg_ses failed: Transport error
  "duplicate_of": [],
Defect 2
This is marked as a duplicate of the previous defect.
  "id": 350018,
  "description": "Starting point:
    [root@s6krw3a ~]# XXXXfirmware --type storage-enclosure enclosure firmware ...
    enclosure 5149-091 78C100G 5342,5342 *5343 s6krw3a-hs enclosure 5149-F48 ...
    [root@s6krw3a ~]#
    Scenario:
    Storage enclosure fw update on 1 Falcon enclosure and 1 6k head unit. Build RC4
    [root@s6krw3a ~]# date; time XXXXfirmware --type storage-enclosure
    Thu May 29 05:26:19 PM EDT 2025
    Thu May 29 05:26:19 PM EDT 2025: XXXXfirmware: Processing node s6krw3a-hs....
    Thu May 29 05:26:23 PM EDT 2025: XXXXfirmware: Processing node s6krw3a-hs....
    s6krw3a-hs: update-directory .
    s6krw3a-hs: [I] Found storage-enclosure firmware update-id firmwareTable ...
    s6krw3a-hs: Updating enclosure firmware ESM_A.
    s6krw3a-hs: [E] Error 1 updating firmware for enclosure sg2.
    inquiry: transport: Host_status=0x0b [DID_SOFT_ERROR] /dev/sg2 doesn’t respond to a SCSI INQUIRY
    sg_ses failed: Transport error
    XXXXfirmware: Command failed. Examine previous error messages to determine cause.
    real 11m56.451s
    user 0m35.035s
    sys 2m7.818s
    [root@s6krw3a ~]#
    But apparently it still worked on canister A, but not on canister B
  "duplicate_of": [
    347852